The perimeter of a triangle is 78cm. The 1st side is 5cm longer than the 2nd side and the 3rd side is 2cm shorter than three times the second. Calculate the length of each side.
Answer by JulietG(1812)

A + B + C = 78
A = B + 5
C = 3B - 2
(B+5) + B + (3B-2) = 78
5B + 3 = 78
5B = 75
B = 15
If B is 15, then C is 43, and A is 20.
15 + 43 + 20 = 78